Find the measure of the central angle of a sector if its area is 5 pi and the radius is 6.
Answer by greenestamps(13215)



If the radius is 6, then the area of the whole circle is 36pi.
The area of the sector is 5pi; the fraction of the circle is (5pi)/(36pi) = 5/36.
A complete circle is 2pi radians; 5/36 of a circle is (5/36)*(2pi) = (5pi)/18 radians.

Answer: the central angle is 5pi/18 radians.